Am totally new to sequences and kinda lost on this problem. Need help for mathletes competition tomorrow
What is the 21st term of the sequence 3,6,12,24

Answer by stanbon(75887)

What is the 21st term of the sequence 3,6,12,24
---------------
1st determine if the sequence is arithmetic or geometric
Yours is geometric with a(1)=3 and r=2
----------------
2nd: Form for the nth term:
a(n) = {a(1))r^(n-1)
---
3rd: a(21) = 3*2^20 = 3* 1048576 = 3145728
==============================================
Cheers,
Stan H.